OpenStudy (anonymous):

medal and fan for best answer!! please help!! What is the missing term, x, in the pattern below? 1, 4, x, 16, 25, … A.12 B.10 C.9 D.8

OpenStudy (accessdenied):

What are some general patterns that you can think of?

OpenStudy (anonymous):

+3, +6, +9 is the only one i can think of

OpenStudy (anonymous):

or more like +3, +6, +6, +9 im not really sure tbh

OpenStudy (accessdenied):

So that is a pattern of adding 3 to each term to get the next. We could also have something where we multiplied by something: 4, x3 = 12, x3 = 36, x3 = 108... This pattern grows much more quickly than one that is additive.

OpenStudy (accessdenied):

Another pattern could be listing all the prime numbers, or perhaps a list of cubed integers (1^3, 2^3, 3^3, 4^3...) Think outside the box a bit, and it might become clear what type of pattern this is!

OpenStudy (anonymous):

i just really cant seem to find what the pattern is.. like it cant be multiplying because then youd have to multiply 1x4but then 16x2 is still bigger than 25 so idk what it would be im so confused!! i dont normally have any problems with these pattern things but this is really difficult!!

OpenStudy (accessdenied):

Have you tried to check the factors of each number? like 4 = 2*2, 16 = 4*4=2*2*2*2, etc. This could also possibly give you a clue.

OpenStudy (anonymous):

i have but 5*5 is the only factors of 25 znd 5 isnt a factor of 16 nor 4

OpenStudy (accessdenied):

25 = 5*5 And 16 = 4*4 this looks suspicious And 4 = 2*2 1 is always able to be multiplied to itself, 1 = 1*1

OpenStudy (accessdenied):

it seems like we have a list of numbers multiplied to themselves, or perfect squares. 1^2, 2^2, ... , 4^2, 5^2

OpenStudy (anonymous):

i got it!! it was +3 +5 +7 +9